# Theoretical and Empirical Insights into the Origins of Degree Bias in Graph Neural Networks

### v1

@@ Abstract
Graph Neural Networks (GNNs) often perform better for high-degree nodes than low-degree nodes on node classification tasks. This degree bias can reinforce social marginalization by, e.g., sidelining authors of lowly-cited papers when predicting paper topics in citation networks. While researchers have proposed numerous hypotheses for why GNN degree bias occurs, we find via a survey of 38 degree bias papers that these hypotheses are often not rigorously validated, and can even be contradictory. Thus, we provide an analysis of the origins of degree bias in message-passing GNNs with different graph filters. We prove that high-degree test nodes tend to have a lower probability of misclassification regardless of how GNNs are trained. Moreover, we show that degree bias arises from a variety of factors that are associated with a node’s degree (e.g., homophily of neighbors, diversity of neighbors). Furthermore, we show that during training, some GNNs may adjust their loss on low-degree nodes more slowly than on high-degree nodes; however, with sufficiently many epochs of training, message-passing GNNs can achieve their maximum possible training accuracy, which is not significantly limited by their expressive power. Throughout our analysis, we connect our findings to previously-proposed hypotheses for the origins of degree bias, supporting and unifying some while drawing doubt to others. We validate our theoretical findings on 8 common real-world networks, and based on our theoretical and empirical insights, describe a roadmap to alleviate degree bias. Our code can be found at: github.com/ArjunSubramonian/degree-bias-exploration.
@@ 1 Introduction
To illustrate this phenomenon, Figure 1 shows that across different message-passing GNNs (cf. §D for details about architectures) applied to the CiteSeer dataset (where nodes represent documents and the classification task is to predict their topic), high-degree nodes generally incur a lower test loss than low-degree nodes. In practice, if such GNNs are applied to predict the topic of documents in social scientific studies, less-cited documents will be misclassified, which can lead to the contributions of their authors not being appropriately recognized and erroneous scientific results. We present additional evidence of degree bias across different GNNs and datasets in §E.
@@
Hence, we theoretically analyze the origins of degree bias during test and training time for general message-passing GNNs, with separate parameters for source and target nodes and residual connections. Our analysis spans different graph filter choices: RW (random walk-normalized filter), SYM (symmetric-normalized filter), and ATT (attention-based filter) (cf. §D for formal definitions). In particular, we prove that high-degree test nodes tend to have a lower probability of misclassification regardless of how GNNs are trained. Moreover, we show that degree bias arises from a variety of factors that are associated with a node’s degree (e.g., homophily of neighbors, diversity of neighbors). Furthermore, we show that during training, SYM (compared to RW) may adjust its loss on low-degree nodes more slowly than on high-degree nodes; however, with sufficiently many epochs of training, message-passing GNNs can achieve their maximum possible training accuracy, which is only trivially curtailed by their expressive power. Throughout our analysis, we connect our findings to previously-proposed hypotheses for the origins of degree bias, supporting and unifying some while drawing doubt to others. We validate our theoretical findings on 8 real-world datasets (cf. §C)that are commonly used in degree bias papers (cf. Figure 2, §F). Based on our theoretical and empirical insights, we describe a principled roadmap to alleviate degree bias.
Figure 1: Test loss vs. degree of nodes in CiteSeer for RW, SYM, and ATT GNNs (over 10 random seeds). High-degree nodes generally incur a lower test loss than low-degree nodes do.
@@ 2 Background and Related Work
We find that the hypotheses are often not rigorously validated, and can even be contradictory. For example, some hypotheses locate the source of degree bias in the training stage, while others cite interactions between training and test-time factors or purely test-time issues. Moreover, hypothesis (H5) in Table 2, which posits that high-degree node representations cluster more strongly, conflicts with and (H10), which argues that high-degree node representations have a larger variance. In our theoretical analysis of the origins of degree bias, we connect our findings to these hypotheses.
We further find that almost no prior works on degree bias provide a comprehensive theoretical analysis of the origins of GNN degree bias that explicitly links a node’s degree to its test and training error in the semi-supervised learning setting (cf. Table 3). For example, most works prove that: (a) high-degree nodes have a larger influence on GNN node representations or parameter gradients, or (b) high-degree nodes cluster more strongly around their class centers or are more likely to be linearly separable; however, these works do not directly bound the probability of misclassifying a node during training vs. test time in terms of its degree.
The few works that do provide a theoretical analysis of degree bias: (A1) perform this analysis with overly strong assumptions, e.g., that graphs are sampled from a Contextual Stochastic Block Model (CSBM) [9], or (A2) posit that GNNs do not have sufficient expressive power to map nodes with different degrees to distinct representations. However, in the case of (1), for CSBM graphs, as the number of nodes $n\to\infty$, the degrees of nodes in each class concentrate around a constant value, making CSBM an inappropriate model to theoretically analyze degree bias. Moreover, many real-world social networks exhibit a power-law degree distribution [2], which is not captured by a CSBM. In the case of (2), §3.4 shows that the accuracy of GNNs on real-world networks is not significantly limited by the Weisfeiler-Leman (WL) test, which draws doubt to hypothesis (H7).
@@
Solutions to degree bias
Numerous papers have contributed solutions to degree bias (cf. §A for a thorough overview). One line of research has produced neighborhood augmentation strategies. For example, [32] perform feature-adaptive neighborhood translation from high-degree nodes to structurally-limited low-degree nodes to enhance their representations. [46] generate multiple views of node neighborhoods (e.g., via node and edge dropping) and learn to maximize the similarity of representations of different views, towards improving the robustness of low-degree node representations. [20] patch the ego-graphs of low-degree nodes by generating virtual neighbors. [49] self-distill graphs and complete the neighborhoods of low-degree nodes with more homophilic links.
Another line of research has produced algorithms that normalize graph filters or node representations to minimize distributional differences between the representations of nodes with different degrees. [21] propose pre-processing and in-processing approaches that re-normalize the graph filter to be doubly stochastic. [50] calculate node representation statistics via a hybrid strategy and use these statistics to normalize representations.
A different line of research has produced algorithms that separate the learning process for high and low-degree nodes. [47] embed degree-specific weights and hashing functions within the layers of GNNs that guarantee degree-aware node representations. [40] propose a degree-specific GNN layer, to avoid parameter sharing across nodes with different degrees, and a Bayesian teacher network that generates pseudo-labeled neighbors for low-degree nodes to increase their proximity to labels.
There exist yet other solutions (e.g., based on adversarial learning, attention) that have been produced. [33] learn debiasing functions that distill and complement the GNN encodings of high and low-degree nodes, respectively. [31] leverage meta-learning to learn to learn representations for low-degree nodes in a locality-aware manner. [17] propose a label proximity score, which they find to be more strongly associated with performance than degree, and learn a new graph structure that reduces discrepancies in label proximity scores across nodes. [57] leverage an attention mechanism to enhance focus on low-degree nodes.
3 Theoretical Analysis
@@
Throughout our theoretical analysis, we connect our findings to previously-proposed hypotheses for the origins of degree bias, supporting and unifying some while drawing doubt to others. We further validate our findings on 8 real-world datasets (cf. §C) that are commonly used in degree bias papers (cf. Figure 2, §F). In all figures (except the PCA plots), error bars are reported over 10 random seeds. The factors of variability include model parameter initialization and training dynamics. All error bars are 1-sigma and represent the standard deviation (not standard error) of the mean. We implicitly assume that errors are normally-distributed. Error bars are computed using PyTorch’s std function [35]. We relegate all proofs to §B.
3.1 Preliminaries
@@
Suppose we have a $C$-class node classification problem defined over an undirected connected graph ${\cal G}=({\cal V},{\cal E})$ with $N=|{\cal V}|$ nodes. We assume that our graph structure ${\bm{A}}\in\{0,1\}^{N\times N}$ and node labels ${\bm{Y}}\in[C]^{N}$ are fixed, but our node features ${\bm{X}}\in\mathbb{R}^{N\times d^{(0)}}$ are independently sampled from class-specific feature distributions, i.e., $\forall i\in{\cal V},{\bm{X}}_{i}\sim{\cal D}_{{\bm{Y}}_{i}}$. We further have a cross-entropy loss function $\ell({\cal M}|i,{\bm{X}},{\bm{A}},c)=\ell({\cal M}|i,c)$ that computes the loss for node $i\in{\cal V}$ with respect to class $c$ for a model ${\cal M}$ that maps ${\bm{X}},{\bm{A}}$ to predictions $\widehat{{\bm{Y}}}\in\mathbb{R}^{N\times C}$. Per the semi-supervised learning paradigm [23], we train ${\cal M}$ with the full graph ${\bm{X}},{\bm{A}}$ but only a labeled subset of nodes $S\subset{\cal V}$.
3.2 Test-Time Degree Bias
In this section, we prove that high-degree test nodes tend to have a lower probability of misclassification. Moreover, we show that GNN degree bias arises from a variety of factors that are associated with a node’s degree (e.g., homophily of neighbors, diversity of neighbors).
@@ 3.2 Test-Time Degree Bias
We first present a theorem that bounds the probability of a test node $i\in{\cal V}\setminus S$ being misclassified. We suppose $\cal M$ is a neural network that has $L$ layers. It takes as input ${\bm{X}},{\bm{A}}$ and generates node representations ${\bm{Z}}^{(L)}\in\mathbb{R}^{N\times C}$; these representations are then passed through a softmax activation function to get $\hat{{\bm{Y}}}={\bm{H}}^{(L)}=\text{softmax}\left({\bm{Z}}^{(L)}\right)$. At this point, we make few assumptions about the architecture of $\cal M$; $\cal M$ could be a graph neural network (GNN), or even an MLP or logistic regression model.
@@
$\mathbb{P}\left(\ell({\cal M}|i,c)>\ell({\cal M}|i,c^{\prime})\right)$ $\leq\frac{1}{1+R_{i,c^{\prime}}},$
@@
Notably, it is not possible to prove a similar lower bound without making assumptions about the higher-order moments of ${\bm{Z}}_{i,c^{\prime}}^{(L)}-{\bm{Z}}_{i,c}^{(L)}$. The coefficient of variation $\frac{\text{Std}\left[{\bm{Z}}_{i,c^{\prime}}^{(L)}-{\bm{Z}}_{i,c}^{(L)}\right]}{\mathbb{E}\left[{\bm{Z}}_{i,c^{\prime}}^{(L)}-{\bm{Z}}_{i,c}^{(L)}\right]}$ is a normalized measure of dispersion that is often used in economics to quantify inequality [12]. Thus, $R_{i,c^{\prime}}$ captures how little $Z_{i}$ varies relative to its expected value. In summary, the probability of misclassification $\mathbb{P}\left(\ell({\cal M}|i,c)>\ell({\cal M}|i,c^{\prime})\right)$ can be minimized when $R_{i,c^{\prime}}$ is maximized. Intuitively, the probability of misclassification is reduced when ${\bm{Z}}_{i}$ is farther away from the $c^{\prime}-c$ decision boundary (i.e., the boundary that separates classes $c$ and $c^{\prime}$) in expectation and has low dispersion. The following subsections reveal why $R_{i,c^{\prime}}$ is large when $i$ is high-degree.
Figure 2: Visual summary of the geometry of representations, variance of representations, and training dynamics of RW, SYM, ATT, and FAIR ATT GNNs on CiteSeer. We consider low-degree nodes to be the 100 nodes with the smallest degrees and high-degree nodes to be the 100 nodes with the largest degrees. Each point in the plots in the left column corresponds to a test node representation and its color represents the node’s class. The plots in the left column are based on a single random seed, while the plots in the middle and right columns are based on 10 random seeds. RW representations of low-degree nodes often have a larger variance than high-degree node representations, while SYM representations of low-degree nodes often have a smaller variance. Furthermore, SYM generally adjusts its training loss on low-degree nodes less rapidly.
3.2.1 Random Walk Graph Filter
@@
${\bm{H}}^{(l)}=\sigma^{(l)}\left({\bm{Z}}^{(l)}\right)=\sigma^{(l)}\left({\bm{H}}^{(l-1)}{\bm{W}}_{1}^{(l)}+{\bm{P}}^{(l)}{\bm{H}}^{(l-1)}{\bm{W}}_{2}^{(l)}+{\bm{X}}{\bm{W}}_{3}^{(l)}\right),$
@@
We first consider the special case that $\forall l\in[L],{\bm{P}}^{(l)}={\bm{P}}_{\text{rw}}={\bm{D}}^{-1}{\bm{A}}$ (i.e., the uniform random walk transition matrix), where ${\bm{D}}$ is the diagonal degree matrix with entries ${\bm{D}}_{ii}=\sum_{j\in{\cal V}}{\bm{A}}_{ij}$. We further simplify the model by choosing all $\sigma^{(l)}$ ($l<L$) to be the identity function (e.g., as in [45]). By doing so, we get the following linear jumping knowledge model $\overline{\text{RW}}$ [52]:
${\bm{H}}^{(L)}=\text{softmax}\left({\bm{Z}}^{(L)}\right)=\text{softmax}\left(\sum_{l=0}^{L}{\bm{P}}^{l}_{\text{rw}}{\bm{X}}{\bm{W}}^{(l)}\right),$
@@
where $\forall l\in[L],{\bm{W}}^{(l)}\in\mathbb{R}^{d^{(0)}\times C}$. For each model ${\cal M}$, $\overline{\cal M}$ denotes the linearized version of the model that we theoretically analyze. Linearizing GNNs is a common practice in the literature [45, 6, 34].
@@
We now prove a lower bound for $R_{i,c^{\prime}}$. By identifying nodes for which this lower bound is larger, we can indirectly figure out which nodes have a lower probability of misclassification. In particular, we find that the bound is generally larger for high-degree nodes, which sheds light on the origins of degree bias. For simplicity of notation, we denote the weights corresponding to the $c^{\prime}-c$ decision boundary (i.e., the boundary that separates classes $c$ and $c^{\prime}$) of the $l$-th term by ${\bm{w}}^{(l)}_{c^{\prime}-c}={\bm{W}}^{(l)}_{.,c^{\prime}}-{\bm{W}}^{(l)}_{.,c}$, and ${\cal N}^{(l)}(i)$ to be the distribution over the terminal nodes of length-$l$ uniform random walks starting from node $i$. We further define:
$\beta_{i,c^{\prime}}^{(l)}=\mathbb{E}_{j\sim{\cal N}^{(l)}(i)}\left[\mathbb{E}_{{\bm{x}}\sim{\cal D}_{{\bm{Y}}_{j}}}\left[{\bm{x}}^{T}{\bm{w}}^{(l)}_{c^{\prime}-c}\right]\right]$
@@
as the $l$-hop prediction homogeneity. In essence, $\mathbb{E}_{{\bm{x}}\sim{\cal D}_{{\bm{Y}}_{j}}}\left[{\bm{x}}^{T}{\bm{w}}^{(l)}_{c^{\prime}-c}\right]$ captures the expected prediction score of ${\bm{w}}^{(l)}_{c^{\prime}-c}$ for a node $j$ whose features ${\bm{X}}_{j}\sim{\cal D}_{{\bm{Y}}_{j}}$; when $\mathbb{E}_{{\bm{x}}\sim{\cal D}_{{\bm{Y}}_{j}}}\left[{\bm{x}}^{T}{\bm{w}}^{(l)}_{c^{\prime}-c}\right]$ is more negative on average, ${\bm{w}}^{(l)}_{c^{\prime}-c}$ predicts $j$ to belong to class $c$ with higher likelihood. Thus, $\beta_{i,c^{\prime}}^{(l)}$ measures the expected prediction score for nodes $j$, weighted by their probability of being reached by a length-$l$ random walk starting from $i$. We also denote the $l$-hop collision probability by:
$\alpha^{(l)}_{i}=\sum_{j\in{\cal V}}\left[\left({\bm{P}}_{\text{rw}}^{l}\right)_{ij}\right]^{2},$
@@
Assume that $\forall l\in[L],\forall j\in{\cal V},\text{Var}_{{\bm{x}}\sim{\cal D}_{{\bm{Y}}_{j}}}\left[{\bm{x}}^{T}{\bm{w}}^{(l)}_{c^{\prime}-c}\right]\leq M$. Then:
$R_{i,c^{\prime}}$ $\geq\frac{\left(\sum_{l=0}^{L}\beta_{i,c^{\prime}}^{(l)}\right)^{2}}{M(L+1)\sum_{l=0}^{L}\alpha^{(l)}_{i}}.$

### v2

@@ Abstract
Graph Neural Networks (GNNs) often perform better for high-degree nodes than low-degree nodes on node classification tasks. This degree bias can reinforce social marginalization by, e.g., privileging celebrities and other high-degree actors in social networks during social and content recommendation. While researchers have proposed numerous hypotheses for why GNN degree bias occurs, we find via a survey of 38 degree bias papers that these hypotheses are often not rigorously validated, and can even be contradictory. Thus, we provide an analysis of the origins of degree bias in message-passing GNNs with different graph filters. We prove that high-degree test nodes tend to have a lower probability of misclassification regardless of how GNNs are trained. Moreover, we show that degree bias arises from a variety of factors that are associated with a node’s degree (e.g., homophily of neighbors, diversity of neighbors). Furthermore, we show that during training, some GNNs may adjust their loss on low-degree nodes more slowly than on high-degree nodes; however, with sufficiently many epochs of training, message-passing GNNs can achieve their maximum possible training accuracy, which is not significantly limited by their expressive power. Throughout our analysis, we connect our findings to previously-proposed hypotheses for the origins of degree bias, supporting and unifying some while drawing doubt to others. We validate our theoretical findings on 8 common real-world networks, and based on our theoretical and empirical insights, describe a roadmap to alleviate degree bias. Our code can be found at: github.com/ArjunSubramonian/degree-bias-exploration.
@@ 1 Introduction
To illustrate this phenomenon, Figure 1 shows that across different message-passing GNNs (see §D for details about architectures) applied to the CiteSeer dataset (where nodes represent documents and the classification task is to predict their topic), high-degree nodes generally incur a lower test loss than low-degree nodes. In practice, if such GNNs are applied to predict the topic of documents in social scientific studies, less-cited documents will be misclassified, which can lead to the contributions of their authors not being appropriately recognized and erroneous scientific results. We present additional evidence of degree bias across different GNNs and datasets in §E.
@@
Hence, we theoretically analyze the origins of degree bias in node classification during test and training time for general message-passing GNNs, with separate parameters for source and target nodes and residual connections. Our analysis spans different graph filter choices: RW (random walk-normalized filter), SYM (symmetric-normalized filter), and ATT (attention-based filter) (see §D for formal definitions). In particular, we prove that high-degree test nodes tend to have a lower probability of misclassification regardless of how GNNs are trained. Moreover, we show that degree bias arises from a variety of factors that are associated with a node’s degree (e.g., homophily of neighbors, diversity of neighbors). Furthermore, we show that during training, SYM (compared to RW) may adjust its loss on low-degree nodes more slowly than on high-degree nodes; however, with sufficiently many epochs of training, message-passing GNNs can achieve their maximum possible training accuracy, which is only trivially curtailed by their expressive power. Throughout our analysis, we connect our findings to previously-proposed hypotheses for the origins of degree bias, supporting and unifying some while drawing doubt to others. We validate our theoretical findings on 8 real-world datasets (see §C)that are commonly used in degree bias papers (see Figure 2, §F). Based on our theoretical and empirical insights, we describe a principled roadmap to alleviate degree bias.
Figure 1: Test loss vs. degree of nodes in CiteSeer for RW, SYM, and ATT GNNs. High-degree nodes generally incur a lower test loss than low-degree nodes do. Error bars are reported over 10 random seeds; all error bars are 1-sigma and represent the standard deviation about the mean.
@@ 2 Background and Related Work
While many of these papers have contributed solutions to degree bias (see §A for a thorough overview), we find that their hypotheses for the origins of degree bias are often not rigorously validated, and can even be contradictory. For example, some hypotheses locate the source of degree bias in the training stage, while others cite interactions between training and test-time factors or purely test-time issues. Moreover, hypothesis (H5) in Table 2, which posits that high-degree node representations cluster more strongly, conflicts with and (H10), which argues that high-degree node representations have a larger variance. In our theoretical analysis of the origins of degree bias, we connect our findings to these hypotheses.
We further find that almost no prior works on degree bias provide a comprehensive theoretical analysis of the origins of GNN degree bias that explicitly links a node’s degree to its test and training error in the semi-supervised learning setting (see Table 3). For example, most works prove that: (a) high-degree nodes have a larger influence on GNN node representations or parameter gradients, or (b) high-degree nodes cluster more strongly around their class centers or are more likely to be linearly separable; however, these works do not directly bound the probability of misclassifying a node during training vs. test time in terms of its degree.
The few works that do provide a theoretical analysis of degree bias: (A1) perform this analysis with overly strong assumptions, e.g., that graphs are sampled from a Contextual Stochastic Block Model (CSBM) [10], or (A2) posit that GNNs do not have sufficient expressive power to map nodes with different degrees to distinct representations. However, in the case of (1), for CSBM graphs, as the number of nodes $n\to\infty$, the degrees of nodes in each class concentrate around a constant value, which is contradictory to real-world graphs, making CSBM an inappropriate model to theoretically analyze degree bias. Moreover, many real-world social networks exhibit a power-law degree distribution [3], which is not captured by a CSBM. In the case of (2), §I shows that the accuracy of GNNs on real-world networks is not significantly limited by the Weisfeiler-Leman (WL) test, which draws doubt to hypothesis (H7).
Ultimately, previously-proposed hypotheses for why GNN degree bias occurs lack rigorous validation, and can even be contradictory. To unify and distill these hypotheses, we provide an analysis of the origins of degree bias in message-passing GNNs with different graph filters.
@@
3 Preliminaries
@@
Throughout our theoretical analysis, we connect our findings to previously-proposed hypotheses for the origins of degree bias, supporting and unifying some while drawing doubt to others. We further validate our findings on 8 real-world datasets (see §C) that are commonly used in degree bias papers (see Figure 2, §F). In all figures (except the PCA plots), error bars are reported over 10 random seeds. The factors of variability include model parameter initialization and training dynamics. All error bars are 1-sigma and represent the standard deviation (not standard error) of the mean. We implicitly assume that errors are normally-distributed. Error bars are computed using PyTorch’s std function [40]. We relegate all proofs to §B.
@@
We first introduce relevant notation and assumptions. Suppose we have a $C$-class node classification problem defined over an undirected connected graph ${\cal G}=({\cal V},{\cal E})$ with $N=|{\cal V}|$ nodes. We assume that our graph structure ${\bm{A}}\in\{0,1\}^{N\times N}$ and node labels ${\bm{Y}}\in\mathbb{N}_{\leq C}^{N}$ are fixed, but our node features ${\bm{X}}\in\mathbb{R}^{N\times d^{(0)}}$ are independently sampled from class-specific feature distributions, i.e., $\forall i\in{\cal V},{\bm{X}}_{i}\sim{\cal D}_{{\bm{Y}}_{i}}$. We further have a model ${\cal M}$ that maps ${\bm{X}},{\bm{A}}$ to predictions $\widehat{{\bm{Y}}}\in\mathbb{R}^{N\times C}$. We use a cross-entropy loss function $\ell({\cal M}|i,c)=-\log\widehat{{\bm{Y}}}_{i,c}$ that computes the loss for node $i\in{\cal V}$ with respect to class $c$ for $\cal M$. Per the semi-supervised learning paradigm [24], we train ${\cal M}$ with the full graph ${\bm{X}},{\bm{A}}$ but only a labeled subset of nodes $S\subset{\cal V}$.
4 Test-Time Degree Bias
@@
The test-time degree bias of models is important to study, as it can yield disparate performance for low-degree nodes when models are deployed in the real world. We prove that high-degree test nodes tend to have a lower probability of misclassification. Moreover, we show that GNN degree bias arises from a variety of factors that are associated with a node’s degree (e.g., homophily of neighbors, diversity of neighbors). We first present a theorem that bounds the probability of a test node $i\in{\cal V}\setminus S$ being misclassified. We suppose $\cal M$ is a neural network that has $L$ layers. It takes as input ${\bm{X}},{\bm{A}}$ and generates node representations ${\bm{Z}}^{(L)}\in\mathbb{R}^{N\times C}$; these representations are then passed through a softmax activation function to get $\hat{{\bm{Y}}}={\bm{H}}^{(L)}=\text{softmax}\left({\bm{Z}}^{(L)}\right)$. At this point, we make few assumptions about the architecture of $\cal M$; $\cal M$ could be a graph neural network (GNN), or even an MLP or logistic regression model.
@@
$\mathbb{P}\left(\ell({\cal M}|i,c)>\ell({\cal M}|i,c^{\prime})\right)$ $\leq\frac{1}{1+R_{i,c^{\prime}}},$ (1)
@@
The assumption that ${\cal M}$ generalizes in expectation is required for the application of Cantelli’s inequality in the proof. Notably, it is not possible to prove a similar lower bound without making assumptions about the higher-order moments of ${\bm{Z}}_{i,c^{\prime}}^{(L)}-{\bm{Z}}_{i,c}^{(L)}$. The coefficient of variation $\frac{\text{Std}\left[{\bm{Z}}_{i,c^{\prime}}^{(L)}-{\bm{Z}}_{i,c}^{(L)}\right]}{\mathbb{E}\left[{\bm{Z}}_{i,c^{\prime}}^{(L)}-{\bm{Z}}_{i,c}^{(L)}\right]}$ is a normalized measure of dispersion that is often used in economics to quantify inequality [13]. Thus, $R_{i,c^{\prime}}$ captures how little $Z_{i}$ varies relative to its expected value. In summary, the probability of misclassification $\mathbb{P}\left(\ell({\cal M}|i,c)>\ell({\cal M}|i,c^{\prime})\right)$ can be minimized when $R_{i,c^{\prime}}$ is maximized. Intuitively, the probability of misclassification is reduced when ${\bm{Z}}_{i}$ is farther away, in expectation, from the decision boundary that separates classes $c$ and $c^{\prime}$, and has low dispersion. The following subsections reveal why $R_{i,c^{\prime}}$ is large when $i$ is high-degree.
Figure 2: Visual summary of the geometry of representations, variance of representations, and training dynamics of RW, SYM, and ATT GNNs on CiteSeer. We consider low-degree nodes to be the 100 nodes with the smallest degrees and high-degree nodes to be the 100 nodes with the largest degrees. Each point in the plots in the left column corresponds to a test node representation and its color represents the node’s class. (In this particular dataset, low-degree nodes are more heavily concentrated in a few classes.) The plots in the left column are based on a single random seed, while the plots in the middle and right columns are based on 10 random seeds. RW representations of low-degree nodes often have a larger variance than high-degree node representations, while SYM representations of low-degree nodes often have a smaller variance. Furthermore, SYM generally adjusts its training loss on low-degree nodes less rapidly.
4.1 Random Walk Graph Filter
@@
${\bm{H}}^{(l)}=\sigma^{(l)}\left({\bm{Z}}^{(l)}\right)=\sigma^{(l)}\left({\bm{H}}^{(l-1)}{\bm{W}}_{1}^{(l)}+{\bm{P}}^{(l)}{\bm{H}}^{(l-1)}{\bm{W}}_{2}^{(l)}+{\bm{X}}{\bm{W}}_{3}^{(l)}\right),$ (2)
@@
We first consider the special case that $\forall l\in\mathbb{N}_{\leq L},{\bm{P}}^{(l)}={\bm{P}}_{\text{rw}}={\bm{D}}^{-1}{\bm{A}}$ (i.e., the uniform random walk transition matrix), where ${\bm{D}}$ is the diagonal degree matrix with entries ${\bm{D}}_{ii}=\sum_{j\in{\cal V}}{\bm{A}}_{ij}$. We further simplify the model by choosing all $\sigma^{(l)}$ ($l<L$) to be the identity function (e.g., as in [51]). By doing so, we get the following linear jumping knowledge model $\overline{\text{RW}}$ [58]:
${\bm{H}}^{(L)}=\text{softmax}\left({\bm{Z}}^{(L)}\right)=\text{softmax}\left(\sum_{l=0}^{L}{\bm{P}}^{l}_{\text{rw}}{\bm{X}}{\bm{W}}^{(l)}\right),$ (3)
@@
where $\forall l\in\mathbb{N}_{\leq L},{\bm{W}}^{(l)}\in\mathbb{R}^{d^{(0)}\times C}$. ${\bm{W}}^{(l)}$ is the sum of all the weight terms that correspond to ${\bm{P}}_{rw}^{l}$ in Eqn. 2; for simplicity, we collapse each sum of weight terms into a single weight matrix. It is still reasonable to have a different weight matrix ${\bm{W}}^{(l)}$ for each term ${\bm{P}}_{rw}^{l}{\bm{X}}$, as we may need to extract different information from features aggregated from neighborhoods at different hops. For each model ${\cal M}$, $\overline{\cal M}$ denotes the linearized version of the model that we theoretically analyze. Linearizing GNNs is a common practice in the literature [51, 7, 39].
@@
We now prove a lower bound for $R_{i,c^{\prime}}$. By identifying nodes for which this lower bound is larger, we can indirectly figure out which nodes have a lower probability of misclassification. In particular, we find that the bound is generally larger for high-degree nodes, which sheds light on the origins of degree bias. For simplicity of notation, we denote the weights corresponding to the decision boundary of the $l$-th term that separates classes $c$ and $c^{\prime}$ by ${\bm{w}}^{(l)}_{c^{\prime}-c}={\bm{W}}^{(l)}_{.,c^{\prime}}-{\bm{W}}^{(l)}_{.,c}$, and ${\cal N}^{(l)}(i)$ to be the distribution over the terminal nodes of length-$l$ uniform random walks starting from node $i$. We further define:
$\beta_{i,c^{\prime}}^{(l)}=\mathbb{E}_{j\sim{\cal N}^{(l)}(i)}\left[\mathbb{E}_{{\bm{x}}\sim{\cal D}_{{\bm{Y}}_{j}}}\left[{\bm{x}}^{T}{\bm{w}}^{(l)}_{c^{\prime}-c}\right]\right]$ (4)
@@
as the $l$-hop prediction homogeneity of $i$ with respect to $c^{\prime}$ when ${\bm{Y}}_{i}=c$. In essence, $\mathbb{E}_{{\bm{x}}\sim{\cal D}_{{\bm{Y}}_{j}}}\left[{\bm{x}}^{T}{\bm{w}}^{(l)}_{c^{\prime}-c}\right]$ captures the expected prediction score of ${\bm{w}}^{(l)}_{c^{\prime}-c}$ for a node $j$ whose features ${\bm{X}}_{j}\sim{\cal D}_{{\bm{Y}}_{j}}$; when $\mathbb{E}_{{\bm{x}}\sim{\cal D}_{{\bm{Y}}_{j}}}\left[{\bm{x}}^{T}{\bm{w}}^{(l)}_{c^{\prime}-c}\right]$ is more negative on average, ${\bm{w}}^{(l)}_{c^{\prime}-c}$ predicts $j$ to belong to class $c$ with higher likelihood. Thus, $\beta_{i,c^{\prime}}^{(l)}$ measures the expected prediction score for nodes $j$, weighted by their probability of being reached by a length-$l$ random walk starting from $i$.
From a topological perspective, because $\beta_{i,c^{\prime}}^{(l)}$ depends on the distribution of random walks from $i$, it is intimately related to local graph structure. Indeed, $\beta_{i,c^{\prime}}^{(l)}$ can be interpreted as a “local subgraph difference” and is highly influenced by the local homophily of $i$. However, $\beta_{i,c^{\prime}}^{(l)}$ is also influenced by the presence of $l$-hop neighbors contained in the training set, as the model is more likely to correctly classify these nodes by a large margin; hence, $\beta_{i,c^{\prime}}^{(l)}$ does not only boil down to local homophily. We discuss other connections between prediction homogeneity, homophily, and separability in §A.4.
In addition to the $l$-hop prediction homogeneity, we denote the $l$-hop collision probability by:
$\alpha^{(l)}_{i}=\sum_{j\in{\cal V}}\left[\left({\bm{P}}_{\text{rw}}^{l}\right)_{ij}\right]^{2},$ (5)
@@
Assume that $\forall l\in\mathbb{N}_{\leq L},\forall j\in{\cal V},\text{Var}_{{\bm{x}}\sim{\cal D}_{{\bm{Y}}_{j}}}\left[{\bm{x}}^{T}{\bm{w}}^{(l)}_{c^{\prime}-c}\right]\leq M$. Then:
$R_{i,c^{\prime}}$ $\geq\frac{\left(\sum_{l=0}^{L}\beta_{i,c^{\prime}}^{(l)}\right)^{2}}{M(L+1)\sum_{l=0}^{L}\alpha^{(l)}_{i}}.$ (6)